OpenStudy (anonymous):

Which expression is equal to sec x° multiplied by cot x°? Answer adjacent ÷ hypotenuse opposite ÷ adjacent hypotenuse ÷ opposite hypotenuse ÷ adjacent

OpenStudy (ash2326):

sec x* cot x sec x =hypoteneus/adjacent cot x =adjacent/ opposite on multiplying hypoteneus/opposite option 3

OpenStudy (asnaseer):

another way of looking at this is as follows:\[\sec(x)=\frac{1}{\cos(x)}\]\[\cot(x)=\frac{1}{\tan(x)}\]and you know:\[\tan(x)=\frac{\sin(x)}{\cos(x)}\]therefore:\[\cot(x)=\frac{1}{\tan(x)}=\frac{\cos(x)}{\sin(x)}\]therefore:\[\sec(x)\times\cot(x)=\frac{1}{\cos(x)}\times\frac{\cos(x)}{\sin(x)}=\frac{1}{\sin(x)}\]\[\sin(x)=\frac{opposite}{hypotenuse}\]so you should be able to guess the answer from here...